# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across various fields, from natural language understanding to text generation. Compared to non-generative LLMs like BERT and DeBERTa, generative LLMs like GPT series and Llama series are currently the main focus due to their superior algorithmic performance. The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. Various hardware platforms exhibit distinct hardware characteristics, which can help improve LLM inference performance. Therefore, this paper comprehensively surveys efficient generative LLM inference on different hardware platforms. First, we provide an overview of the algorithm architecture of mainstream generative LLMs and delve into the inference process. Then, we summarize different optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP, and provide inference results for generative LLMs. Furthermore, we perform a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms by considering hardware power consumption, absolute inference speed (tokens/s), and energy efficiency (tokens/J). We compare the performance of the same optimization methods across different hardware platforms, the performance across different hardware platforms, and the performance of different methods on the same hardware platform. This provides a systematic and comprehensive summary of existing inference acceleration work by integrating software optimization methods and hardware platforms, which can point to the future trends and potential developments of generative LLMs and hardware technology for edge-side scenarios. Our project is available at https://dai.sjtu.edu.cn/project.html.
Keywords Generative Large Language Model  $\cdot$ Hardware  $\cdot$ CPU  $\cdot$ GPU  $\cdot$ FPGA  $\cdot$ ASIC  $\cdot$ PIM  $\cdot$ NDP
1 Introduction
@@
Large Language Models (LLMs) have become cornerstones of modern artificial intelligence, demonstrating remarkable capabilities across a spectrum of fields, from natural language understanding to text generation [1, 2, 3, 4, 5]. LLMs can be categorized into two primary types: generative LLMs and non-generative LLMs. Non-generative LLMs, such as BERT [6], RoBERTa [7], ELECTRA [8], and DeBERTa [9], are designed to classify and make predictions based on input text. These models typically range in size from millions of parameters, allowing them to excel in tasks that require discernment and nuanced understanding. BERT, introduced in 2018, only has 340 million parameters. RoBERTa, introduced in 2019, slightly increases to 355 million parameters. And DeBERTa, released in 2021, increases to 1.5 billion. Generative LLMs, like GPT series [10, 11, 12, 13], T5 [14], OPT [15], BLOOM [16], and Llama series [17, 18, 19], have taken language generation to new heights. The model size increase of generative LLMs [10, 11, 20, 12, 14, 21, 22, 23, 24, 25, 26, 27, 28, 29, 15, 16, 30, 13, 31, 32, 33, 34, 35, 17, 18, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 19, 51, 52, 53, 54, 55, 56, 57] are particularly notable in the past 6 years, as shown in Figure 1. In 2018, GPT1 has only 110 million parameters, which grows to 1.5 billion in GPT2 in 2019. GPT3, launched in 2020, grows to 175B parameters dramatically, and GPT3.5 maintains the same size. After 2022, the model size maintains to several hundreds and thousands of billions like GPT4, Llama3, and Grok1 [58]. The evolution of LLMs has been characterized by an exponential growth in model parameters, which has been instrumental in enhancing their performance and versatility. Compared to non-generative LLMs, generative LLMs are currently the primary focus of research and development in the field of LLMs for their superior algorithmic performance. In recent years, after reaching the trillion-parameter scale in 2022, the parameter size of generative LLMs has stopped growing at an exponential rate. Two main reasons can explain this phenomenon: (1) As the amount of computation increases, the demand for computing power also rises significantly. The slow growth of hardware capabilities, particularly the slowing down of Moore’s Law [59], limits the improvement of single-chip computing power. (2) Researchers have found that model performance is not solely dependent on the number of parameters, but also on the quantity and quality of training data [60]. By providing more qualified training tokens, the algorithmic performance can be further improved [18, 19]. At the same time, the parameter sizes of generative LLMs have shifted from "small to large" to "remaining stable" or even "shrinking". More models with fewer parameters are being released, particularly those that are better suited for deployment on edge devices. Notably, OpenAI’s recent o1 [61] improves algorithm performance by introducing Chain-of-Thought (CoT) reasoning and multi-step inference. This new computational paradigm increases the importance of inference within the model, further highlighting the need to accelerate inference efficiency.
Figure 1: Typical LLM model size in the past six years.
@@
The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. Due to the continuation of Moore’s Law, from 2018 to 2022, GPU manufacturing processes have progressed from 12nm to 3nm, and the floating-point performance of single GPU die has increased from 130 TFLOPS to 989 TFLOPS. During model training, GPUs are used predominantly due to the user-friendliness of the CUDA programming stack [62] and the high scalability of GPU chips (e.g. NVLink [63]). During inference, various hardware options like CPU, GPU, FPGA, and ASIC exhibit distinct hardware characteristics, which can help improving LLM inference performance. CPUs offer high programmability with a computing power of approximately 4 to 70 TOPS and with power consumption around from 4W to >200W. Modern CPUs (including some System-on-Chips, SoCs) enhance AI performance by integrating domains-specific architecture (DSA) units. These include Apple’s Neural Engine in the M2 Ultra [64], Qualcomm’s NPU in the Snapdragon 8 Gen3 [65], and Intel’s AVX/AMX ISA extensions [66]. GPUs excel in parallelism and computing power, delivering between $\sim$70 to >1000 TOPS and featuring an impressive memory bandwidth of up to 1555 GB/s. On one hand, GPUs integrate a large number of SIMD cores and Tensor Cores in NVIDIA V100/A100/H100 [67, 68, 69] or Matrix Cores in AMD Instinct MI100/MI200/MI300 series [70, 71, 72] to enhance computing powers. On the other hand, GPUs support lower precision computations, such as INT8, FP8 and INT4 [68, 69], which allows for more multiplication units to be packed into a given chip area. Nevertheless, their power consumption is significantly higher, ranging from $\sim$20W to >700W. FPGAs offer substantial parallelism and optimization capabilities, with computing performance between 50 to 100 TOPS. They are also more power-efficient, consuming about 75 to 100W. ASICs are often designed for specific applications and custom silicon designs, resulting in a wide range of computing power from GOPS to TOPS. Their power consumption can vary from the 0.1W range to several hundred watts. Due to their specialized design, ASICs generally offer higher computational efficiency and better energy efficiency compared to GPUs and CPUs. The unique attributes of each hardware type influence their optimal application in various inference scenarios.
@@
Here, we list and compare the existing surveys of LLM inference in Table 1. Previous surveys [73, 74, 75, 76, 77] primarily summarize various software optimization methods like quantization, sparsity, fast decoding for generative LLMs from an algorithm perspective. However, they do not take into account whether different optimization methods exhibit varying inference performance across different hardware platforms, and similarly, they also lack a fair and quantitative comparison. Surveys [78, 79] focus on accelerating transformer-based LLMs, including non-generative LLMs like BERT and a very small number of generative LLMs like GPT, but merely list the work done on different hardware platforms. They lack a summary and abstraction of the optimization methods used by different accelerators. Additionally, it only provides a relative comparison of speedup and energy efficiency with different baselines while lacking a fair comparison of inference performance, as what matters most for generative LLMs is the absolute inference speed (tokens per second, tokens/s) and inference energy efficiency (tokens per joule, tokens/J). Like [78, 79], surveys [80, 81] mainly focus on non-generative LLMs and one or two specific hardware platforms. Our survey focuses solely on generative LLMs, summarizing various software optimization methods in conjunction with multiple hardware platforms, including CPUs, GPUs, FPGAs, ASICs, and PIM/NDPs. For the first time, we innovatively use performance metrics that matter to generative LLMs: the number of tokens generated per second (tokens/s) and the number of tokens generated per joule (tokens/J). We compare (1) the performance of the same optimization methods across different hardware platforms, (2) the performance across different hardware platforms, and (3) the performance of different methods on the same hardware platform. This provides a systematic and comprehensive summary of existing inference acceleration work by integrating software optimization methods and hardware platforms.
Table 1: Comparison of existing LLM surveys
@@
[73, 74, 75, 76, 77]
[78, 79]
[80]
[81]
@@
The article is meticulously structured to comprehensively summarize different optimizations on different hardware platforms for generative LLMs. Section 2 delves into the inference process of LLMs, providing an overview of the architecture and functioning of mainstream generative LLMs. Section 3 first summarizes the different optimization methods on various platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP in tabular form, and then provides a detailed description of each method and related works. Additionally, for each method, we also perform a qualified and quantitative comparison to show the difference among the hardware platforms. Furthermore, section 4 performs a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms. And we also point to the future trends and potential developments of generative LLMs and hardware technology for edge-side scenarios. Section 5 summarizes the work of this survey.
2 Generative LLM Architecture
2.1 Vanilla Generative LLM
The most common generative LLM is based on the transformer structure due to its abilities for capturing long-term dependencies [82]. The inference of generative LLM consists of two stages, the prefill stage and the decode stage, as shown in Figure 2. In the prefill stage, the input text is converted into embeddings and input into the LLM all at once. Each transformer layer performs operations, and the intermediate calculation results (K and V) are saved in the KV cache of the self-attention block for use in the decode stage. After completing calculating the last layer and LM head, the first token is generated. Then, in the decode stage, the LLM generates each token output autoregressively and updates the KV cache each time until generating the stop string. In the prefill stage, the attention computation has quadratical computational and storage complexity related with input text length. To address this limitation, many hardware-efficient LLMs focus on more efficient processing of longer length. For daily application scenarios where the number of input tokens is from 128 to 256 and output tokens larger than 32, we obtain that the time proportion of the decode stage exceeds 80% by profiling Llama2-7B on single NVIDIA A100 GPU, as shown in Figure 2 (c). Therefore, we summarize these LLMs in sub-section 2.2. And for the decode stage, we summarize hardware optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP, and provide inference results for generative LLMs in latter sub-sections.
@@ 2.1 Vanilla Generative LLM
2.2 Efficient Generative LLM
2.2.1 Transformer-based LLM
@@
Transformer-XL [83] adopts a segment-level recurrence mechanism and a novel positional encoding scheme to learn dependencies beyond a fixed length without disrupting temporal coherence. Linear Transformer [84] represents self-attention as a linear dot product of kernel feature maps and alters the computation order by leveraging the associativity of matrix multiplication. This modification reduces the complexity from $O(L^{2})$ to $O(L)$, where $L$ is the context length, significantly accelerating the computation of autoregressive Transformers. Another efficient structure is the Attention-Free Transformer (AFT) [85]. Unlike vanilla transformers, which first compute the query-key product, AFT combines the key and value with a set of learned positional biases before performing element-wise multiplication with the query. As a result, the memory complexity of AFT is linear with respect to both the context size and feature dimensions, enabling support for larger input lengths and model sizes. Based on AFT, the Receptance Weighted Key Value (RWKV) [86] combines the efficient parallel training capabilities of Transformers with the efficient inference of RNNs. It leverages linear attention mechanisms and allows the model to be expressed as either a transformer or an RNN. It also enables parallel computation during training while maintaining constant computational and memory complexity during inference. DiJiang [87] introduces a novel frequency-domain kernelization method based on the Discrete Cosine Transform (DCT). It points out that improving attention mechanisms often requires extensive retraining, which is impractical for large language models with vast numbers of parameters. This approach enables the conversion of a pre-trained standard Transformer into a model with linear complexity and low training costs, utilizing a weighted quasi-Monte Carlo method for sampling. Extensive experiments demonstrate that this method achieves performance comparable to the vanilla transformer while significantly reducing training costs and substantially increasing inference speed.
2.2.2 SSM-based LLM

### v3

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable capabilities across various fields, from natural language understanding to text generation. Compared to non-generative LLMs like BERT and DeBERTa, generative LLMs like GPT series and Llama series are currently the main focus due to their superior algorithmic performance. The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. Various hardware platforms exhibit distinct hardware characteristics, which can help improve LLM inference performance. Therefore, this paper comprehensively surveys efficient generative LLM inference on different hardware platforms. First, we provide an overview of the algorithm architecture of mainstream generative LLMs and delve into the inference process. Then, we summarize different optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP, and provide inference results for generative LLMs. Furthermore, we perform a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms by considering hardware power consumption, absolute inference speed (tokens/s), and energy efficiency (tokens/J). We compare the performance of the same optimization methods across different hardware platforms, the performance across different hardware platforms, and the performance of different methods on the same hardware platform. This provides a systematic and comprehensive summary of existing inference acceleration work by integrating software optimization methods and hardware platforms. We point out that the development of edge intelligence has gained significant momentum, driven by the increasing capability of LLMs and the increasing demands of edge applications. And three trends (multimodality, inference-time compute, and higher inference energy efficiency) are promising to redefine the capabilities of edge artificial intelligence systems. Our project is available at https://dai.sjtu.edu.cn/project.html.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
Large Language Models (LLMs) have become cornerstones of modern artificial intelligence, demonstrating remarkable capabilities across a spectrum of fields, from natural language understanding to text generation (mo2024large, ; wu2024large, ; li2024pre, ; shu2024rewritelm, ; thakur2024verigen, ). LLMs can be categorized into two primary types: generative LLMs and non-generative LLMs. Non-generative LLMs, such as BERT (devlin2018bert, ), RoBERTa (liu2019roberta, ), ELECTRA (clark2020electra, ), and DeBERTa (he2020deberta, ), are designed to classify and make predictions based on input text. These models typically range in size from millions of parameters, allowing them to excel in tasks that require discernment and nuanced understanding. BERT, introduced in 2018, only has 340 million parameters. RoBERTa, introduced in 2019, slightly increases to 355 million parameters. And DeBERTa, released in 2021, increases to 1.5 billion parameters. Generative LLMs, like GPT series (radford2018improving, ; radford2019language, ; brown2020language, ; achiam2023gpt, ), T5 (raffel2020exploring, ), OPT (zhang2022opt, ), BLOOM (le2023bloom, ), and Llama series (llama, ; llama2, ; meta2024llama3, ), have taken language generation to new heights. The model size increase of generative LLMs (radford2018improving, ; radford2019language, ; lewis2019bart, ; brown2020language, ; raffel2020exploring, ; du2021glm, ; xue2020mt5, ; fedus2021switch, ; zeng2021pangu, ; sun2021ernie, ; sanh2021multitask, ; du2022glam, ; wang2021ernie, ; chowdhery2022palm, ; zhang2022opt, ; le2023bloom, ; hoffmann2022training, ; achiam2023gpt, ; thoppilan2022lamda, ; anthropic2024claude, ; chatglm, ; chatglm2, ; chatglm3, ; llama, ; llama2, ; zheng2023judging, ; almazrouei2023falcon, ; jiang2024mixtral, ; peng2023rwkv, ; xai2024grok, ; jiang2023mistral, ; team2023gemini, ; mamba, ; mamba2, ; team2024gemma, ; openai2024gpt4o, ; bai2023qwen, ; qwen2.5, ; yang2024qwen2, ; qwen1.5, ; meta2024llama3, ; meta2024llama31, ; meta2024llama32, ; reid2024gemini, ; deepmind2024gemini, ; anthropic2024claude3, ; openai2024gpt35turbo, ; grok2024, ) are particularly notable in the past 6 years, as shown in Figure 1. In 2018, GPT1 has only 110 million parameters, which grows to 1.5 billion in GPT2 in 2019. GPT3, launched in 2020, grows to 175B parameters dramatically, and GPT3.5 maintains the same size. After 2022, the model size maintains to several hundreds and thousands of billions like GPT4, Llama3, and Grok1 (grok1, ). The evolution of LLMs is characterized by an exponential growth in model parameters, which enhances their performance and versatility. Compared to non-generative LLMs, generative LLMs are currently the primary focus of research and development in the field of LLMs for their superior algorithmic performance. In recent years, after reaching the trillion-parameter scale in 2022, the parameter size of generative LLMs has stopped growing at an exponential rate. Two main reasons can explain this phenomenon: (1) As the amount of computation increases, the demand for computing power also rises significantly. The slow growth of hardware capabilities, particularly the slowing down of Moore’s Law (schaller1997moore, ), limits the improvement of single-chip computing power. (2) Researchers have found that model performance is not solely dependent on the number of parameters, but also on the quantity and quality of training data (kaplan2020scaling, ). By providing more qualified training tokens, the algorithmic performance can be further improved (llama2, ; meta2024llama3, ). At the same time, the model size of generative LLMs have shifted from "increasing" to "remaining stable" or even "shrinking". More models with fewer parameters are released, which are more suitable for deployment on edge devices.
Figure 1. Typical LLM model size in the past six years.
@@
The advancements in generative LLMs are closely intertwined with the development of hardware capabilities. Due to the continuation of Moore’s Law, from 2018 to 2022, GPU manufacturing processes have progressed from 12nm to 3nm, and the floating-point performance of single GPU die has increased from 130 TFLOPS to 989 TFLOPS. During model training, GPUs are used predominantly due to the user-friendliness of the CUDA programming stack (cook2012cuda, ) and the high scalability of GPU chips (e.g. NVLink (foley2017ultra, )). During inference, various hardware options like CPU, GPU, FPGA, and ASIC exhibit distinct hardware characteristics, which can help improving LLM inference performance. CPUs offer high programmability with a computing power of approximately 4 to 70 TOPS and with power consumption around from 4W to >200W. Modern CPUs (including some System-on-Chips, SoCs) enhance AI performance by integrating domain-specific architecture (DSA) units. These include Apple’s Neural Engine in the M2 Ultra (Apple_M2_Ultra, ), Qualcomm’s NPU in the Snapdragon 8 Gen3 (Snapdragon8Gen3, ), and Intel’s AVX/AMX ISA extensions (4th_Gen_Intel_Xeon, ). GPUs excel in parallelism and computing power, delivering between $\sim$70 to >1000 TOPS and featuring an impressive memory bandwidth of up to 1555 GB/s. On one hand, GPUs integrate a large number of SIMD cores and Tensor Cores in NVIDIA V100/A100/H100 (nvv100, ; nva100, ; nvh100, ) or Matrix Cores in AMD Instinct MI100/MI200/MI300 series (amdcdna1, ; amdcdna2, ; amdcdna3, ) to enhance computing powers. On the other hand, GPUs support lower precision computations, such as INT8, FP8 and INT4 (nva100, ; nvh100, ), which allows for more multiplication units to be packed into a given chip area. Nevertheless, their power consumption is significantly higher, ranging from $\sim$20W to >700W. FPGAs offer substantial parallelism and optimization capabilities, with computing performance between 50 to 100 TOPS. They are also more power-efficient, consuming about 75 to 100W. AMD’s Xilinx offers Zynq (zynq, ), Virtex (virtex, ), and Versal (versal, ) FPGA series for edge-side and cloud-side computing. The Zynq series, like ZCU102, combine ARM processors with FPGA programmable logic, has a power consumption of 10W-30W. The Virtex series, such as the VU13P, are designed for high-performance computing and data centers, providing up to 6.2 TFLOPS of computing power, with 3,840 DSP units and power consumption ranging from 30W to 70W. The Versal series, such as the VCK5000, integrate AI Engines (versal_aie, ) and efficient parallel computation units, offering >50 TOPS of computing power and 4,000 DSP units, with power consumption of 225W. ASICs like Groq LPU (groq_lpu, ) and Cerebras WSE-3 (lie2024wafer, ) are often designed for specific applications with the customized architecture, offering higher computational efficiency and better energy efficiency.
@@
Here, we list and compare the existing LLM inference surveys in Table 1. Previous surveys (miao2023towards, ; zhu2023survey, ; qu2024mobile, ; park2024comprehensive, ; zhou2024survey, ) primarily summarize various software optimization methods like quantization, sparsity, fast decoding for generative LLMs from an algorithm perspective. However, they do not consider that different optimization methods exhibit different inference performance across different hardware platforms, and similarly, they also lack a fair and quantitative comparison. Surveys (kachris2024survey, ; koilia2024hardware, ) focus on accelerating transformer-based LLMs, including non-generative LLMs like BERT and a small number of generative LLMs like GPT, but merely list the work done on different hardware platforms. And they lack a summary and abstraction of the optimization methods used by different accelerators. Additionally, it only provides a relative comparison of speedup and energy efficiency with different baselines, which is unfair. Like (kachris2024survey, ; koilia2024hardware, ), surveys (wolters2024memory, ; kang2024survey, ) mainly focus on non-generative LLMs on one or two specific hardware platforms. Our survey focuses solely on generative LLMs, summarizing various software optimization methods in conjunction with multiple hardware platforms, including CPUs, GPUs, FPGAs, ASICs, and PIM/NDPs. What matters most for generative LLMs is the absolute inference speed and the inference energy efficiency. Therefore, for the first time, we use the number of tokens generated per second (tokens per second, tokens/s) and the number of tokens generated per joule (tokens per joule, tokens/J) to evaluate LLM acceleration. Besides these two metrics, we also conduct the comparison on (1) the performance of the same optimization methods across different hardware platforms, (2) the performance across different hardware platforms, and (3) the performance of different methods on the same hardware platform. This can provide the systematic and comprehensive summary of existing inference accelerations with software optimization methods and hardware platforms.
Table 1. Comparison of existing LLM surveys
@@
(miao2023towards, ; zhu2023survey, ; qu2024mobile, ; park2024comprehensive, ; zhou2024survey, )
(kachris2024survey, ; koilia2024hardware, )
(wolters2024memory, )
(kang2024survey, )
@@
This survey aims to systematically summarize the optimization methods for generative LLMs across different hardware platforms. Section 2 delves into the inference process of LLMs, providing an overview of the architecture and functioning of mainstream generative LLMs. Section 3 first summarizes the different optimization methods on various platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP in tabular form, and then provides a detailed description of each method and related works. Additionally, for each method, we also perform a qualified and quantitative comparison to show the difference among the hardware platforms. Section 4 performs a qualitative and quantitative comparison of inference performance with batch sizes 1 and 8 on different hardware platforms. Furthermore, in section 5, we point out that three trends (multimodality, inference-time compute, and higher inference energy efficiency) are promising to redefine the capabilities of edge artificial intelligence systems. Section 6 summarizes the work of this survey.
2. Generative LLM Architecture
2.1. Overview
The most common generative LLM is based on the transformer structure due to its abilities for capturing long-term dependencies (vaswani2017attention, ). The inference of generative LLM consists of two stages, the prefill stage and the decode stage, as shown in Figure 2. In the prefill stage, the input text is converted into embeddings and input into the LLM all at once. Each transformer layer performs operations, and the intermediate results (Key and Value, KV cache) are saved for reuse during the decode stage. After completing calculating the last layer and LM head, the first token (‘My’) is generated. Then, in the decode stage, the LLM generates each token output autoregressively and updates the KV cache each time until generating the stop string (‘</s>’). For daily application scenarios where the number of input tokens is from 128 to 256 and output tokens larger than 32, we obtain that the time proportion of the decode stage exceeds 80% by profiling Llama2-7B on single NVIDIA A100 GPU, as shown in Figure 2(c).
In the prefill stage, the attention computation has quadratical computational and storage complexity related with input text length. To address this issue, many hardware-efficient LLMs are designed for more efficient processing of longer length. Therefore, we summarize these LLMs in sub-section 2.2. And for the decode stage, we summarize hardware optimization methods for different platforms such as CPU, GPU, FPGA, ASIC, and PIM/NDP, and provide inference results for generative LLMs in latter sub-sections.
@@ 2.1. Overview
2.2. Efficient Generative LLM
2.2.1. Attention-based LLM
In LLMs, the most common type of attention is self-attention, where each token in the input sequence attends to every other token, capturing both local and global dependencies. This is achieved by computing three vectors for each token: Query ($Q$), Key ($K$), and Value ($V$). The attention scores are calculated as the scaled dot product of $Q$ and $K$, followed by a softmax operation to determine the weight of each token in the sequence. These weights are then used to compute a weighted sum of the $V$ vectors, producing the attention output.
(1) $O=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$
@@ 2.2.1. Attention-based LLM
Many methods focus on simplifying the attention mechanism. Transformer-XL (transformer-xl, ) adopts a segment-level recurrence mechanism and a positional encoding scheme to learn dependencies beyond a fixed length without disrupting temporal coherence. Linear Transformer (linear-transformer, ) represents self-attention as a linear dot product of kernel feature maps and alters the computation order by leveraging the associativity of matrix multiplication. This modification reduces the complexity from $O(L^{2})$ to $O(L)$, where $L$ is the context length, significantly accelerating the computation of autoregressive Transformers. Another efficient structure is the Attention-Free Transformer (AFT) (attention-free, ). Unlike vanilla transformers, which first compute the query-key product, AFT combines the key and value with a set of learned positional biases before performing element-wise multiplication with the query. As a result, the memory complexity of AFT is linear with respect to both the context size and feature dimensions, enabling support for larger input lengths and model sizes. Based on AFT, the Receptance Weighted Key Value (RWKV) (peng2023rwkv, ) combines the efficient parallel training capabilities of Transformers with the efficient inference of RNNs. It leverages linear attention mechanisms and allows the model to be expressed as either a transformer or an RNN. It also enables parallel computation during training while maintaining constant computational and memory complexity during inference. DiJiang (dijiang, ) introduces a novel frequency-domain kernelization method based on the Discrete Cosine Transform (DCT). It points out that improving attention mechanisms often requires extensive retraining, which is impractical for large language models with vast numbers of parameters. This approach enables the conversion of a pre-trained standard Transformer into a model with linear complexity and low training costs, utilizing a weighted quasi-Monte Carlo method for sampling. Extensive experiments demonstrate that this method achieves performance comparable to the vanilla transformer while significantly reducing training costs and substantially increasing inference speed.
2.2.2. SSM-based LLM